# TK-Planes: Tiered K-Planes with High Dimensional Feature Vectors for Dynamic UAV-based Scenes

### v1

@@
In this paper, we present a new approach to bridge the domain gap between synthetic and real-world data for unmanned aerial vehicle (UAV)-based perception. Our formulation is designed for dynamic scenes, consisting of moving objects or human actions, where the goal is to recognize the pose or actions. We propose an extension of K-Planes Neural Radiance Field (NeRF), wherein our algorithm stores a set of tiered feature vectors. The tiered feature vectors are generated to effectively model conceptual information about a scene as well as an image decoder that transforms output feature maps into RGB images. Our technique leverages the information amongst both static and dynamic objects within a scene and is able to capture salient scene attributes of high altitude videos. We evaluate its performance on challenging datasets, including Okutama Action and UG2, and observe considerable improvement in accuracy over state of the art aerial perception algorithms.
@@ I Introduction
We have evaluated the performance on some challenging UAV datasets, including Okutama Action [3] and UG2 [39], which consist of high altitude videos. We show that our NeRF-based model based on TK-planes can generate supplementing UAV-based data that can improve the overall recognition accuracy for dynamic scenes, as compared to state of the art algorithms [9, 21]. To summarize, some novel contributions from our work include:
  • Development of a novel grid-based NeRF architecture that utilizes multiple scales of feature space in order to render high fidelity regions of a scene.
  • Extension of previous pipelines for training grid-based NeRF architectures to include an image decoder and address edge effects from rendering image patches.
  • Showcasing high fidelity renderings over previous methods such as K-Planes [9] and Extended K-Planes [21].
@@ II Related Works
Neural Rendering for Dynamic Scenes. To represent 3D dynamic scenes, various NeRF research explores 4D neural radiance fields representation with an additional temporal dimension [6, 43, 10, 40]. Instead of embedding continuous time dimension into the radiance fields, some of variants uses time-varying latent code [18, 26]. While these 4D NeRF show capability to represent complicated dynamic scenes, they generally suffer from inherent representational redundancy when they learn 4D contents from 2D imagery only. The other major variants of NeRF incorporates canonical scene and a deformation field [25, 29]. However, these methods demonstrate limited capability to learn topological changes or large motion in the scene due to their heavy dependence on the canonical scene.
@@ III-A Novel View Synthesis for Dynamic Scenes
Given a difficult domain such as UAV video with limited available training data and a small dynamic to static pixel ratio, our approach utilizes NeRF in feature space in order to better capture and render objects of interest within a scene, e.g. dynamic objects. We employ a grid-based Neural Radiance Field, the grids of which are similar to those from K-Planes [9]. A key difference being that the grids in our algorithm store feature vectors that do not directly code for RGB values but rather for more abstract high level concepts within the scene, such as patches of ground, trees, and people. Thus, our NeRF algorithm outputs a feature vector and density value for a given ray rather than an RGB and density value. The overall output is a feature map rather than an RGB image. Output feature maps are decoded by a convolutional neural network in order to generate a final rendered image.
III-B Feature Plane NeRF Overview
To understand our algorithm in more detail, a brief overview of Neural Radiance Fields follows. Neural radiance fields (NeRF) are a type of volumetric rendering that utilize neural networks to output a color and density given an input camera pose. From a designated camera pose, rays are cast into a scene and incremental points along each ray are sampled. Typically, two Multi-layer Perceptrons (MLP) are employed, one for spatial density and the other for a pixel RGB value. Each sampled point’s location and direction are fed as input into the MLPs. The first MLP takes as input a position within 3D space and outputs a density value as well as a feature vector. The feature vector and the point’s direction are fed into the second MLP which outputs a RGB color value. The density and color at each sampled point along the ray are then multiplied, rectified and summed along the ray to get a final pixel color value for the given ray. This version of NeRF relies on the MLPs learning an implicit 3D structure of a scene.
Alternative to the implicit method, some NeRF algorithms implement an explicit representation of a scene stored in a data structure such as a matrix [15, 38] or a grid [9, 4]. These data structures contain feature vectors that are input into the spatial density MLP and learn explicit details wherein the data structure itself represents the spatial or temporal layout of a scene. By storing feature vectors in an explicit data structure, the MLPs used to output density and color values can be smaller and lightweight which coincides with reduced training and inference times.
@@ III-C Tiered Planes in Feature Space
Our main approach operates in a conceptual feature space rather than an RGB space with respect to the output of the NeRF algorithm. The purpose of focusing on feature space rather than RGB space is to learn feature vectors that represent cohesive objects of a scene, such as people, as opposed to learning solely a chromaticity and luminance value. This helps to alleviate the issue of poor fidelity of high interest objects across different poses and time-slices. We use the same grid representation as seen by K-Planes [9] with changes influenced from the extended version as seen in [21]. Figure 1 shows our system in full with varying scale feature space grids as well as the image decoder.
In particular, we add the additional planes for spatial features that are dynamic as well as the cosine similiarity loss function that encourages separation of static and dynamic features in the spatial planes. The resulting static and dynamic features are described as follows:
@@
Figure 1: Overview of our tiered planes algorithm: Three sets of grids are shown, representing three different scales of feature space, with larger scales capturing more abstract scene details. Note that the algorithm uses nine grids in total per scale, three static spatial, three dynamic spatial and three dynamic spatio-temporal. Feature vectors are processed via concurrent MLPs to output a density value and final feature vector. These are accumulated with volumetric rendering into a feature map. Feature maps from each set of grids are input into subsequent stages of the image decoder.
wherein $f_{s}$ is a function of the static spatial planes and $f_{d}$ is a function of the dynamic spatial and temporal planes and $\prod$ represents the Hadamard product. In our version, each feature vector $f_{*}(\mathbf{q})_{c}$ is a $D$ dimensional vector wherein $D$ is a hyperparameter, which differs from [21] in that it does not include a learned mask value for the temporal feature vectors. In focusing on rendering quality over data generation, we can omit the mask dimension to simplify the model. From 2, a cohesive feature vector $f(\mathbf{q})$ is calculated by concatenating $f_{s}$ and $f_{d}$. $f(\mathbf{q})$ is then input into the MLP networks to output a density value and another feature vector. Volumetric rendering along the ray returns a final feature vector value per pixel of an output feature map at a given feature space scale, $d_{s}$. Furthermore, each set of grids reduces the dimensionality of their feature vectors by two compared to the previous grids. In our experiments we choose a starting $D$ of 256 such that vectors in the first stage of grids can learn complex and abstract details of a scene. The feature map from these grids is fed into the start of the image decoder. Subsequent grid sets then have feature vectors of size 128 and 64 and are fed into later stages of the dcoder, as explained below. The purpose of this is to focus on more fine grained detail in the later grids which does not require large vector lengths. This also allows to grid feature map outputs to integrate with an image decoder in which the spatial dimensions are doubled each stage, but the depth dimension is halved.
K-Planes [9] and its extended version [21] use multi resolution grids offering different levels of scene details. The resulting feature vector from each level can either be summed or concatenated before being processed by the MLPs and then accumulated via volumetric rendering. In these cases, the volumetric rendering is the final step. However, in our algorithm it is more intuitive to separate the different levels of detail as input into appropriate stages of the decoder rather than as input entirely into the front of the decoder. Convolutional image decoders operate on high-level abstract features to low-level detail features as their stages progress and typically benefit from feature map input at varying stages [32]. This allows the grids at varying scales of feature space to learn appropriate details with respect to rendering an output image.
@@
Figure 2: a.) A diagram of a block within the image decoder. The $n^{th}$ block accepts two inputs, a feature map from the $n^{th}$ set of grids after volumetric rendering and the output $O_{n-1}$ from the previous block. At $n=0$, there is no input from a previous block, only the feature map. b.) The final image decoder block with $xN$ 3x3 convolutional layers followed by $xM$ 1x1 convolutional layers and a sigmoid activation layer to rectify the raw output into RGB values.
III-D Image Decoder
The set of feature maps $d_{s}$ at varying scales, $d$, are input to an image decoder that consists of convolutional blocks followed by bilinear upsampling. Figure 2 details a block of the image decoder. Each block takes as input a corresponding feature map and, if applicable, the output of the previous block. These two inputs are concatenated together before further processing. The output of each block has spatial dimensions upsampled by two and feature dimensions downsampled by two. The number of blocks is equivalent to the number of scaling levels of the grids, in our case three. A final block consisting of four convolutional layers further processes the resultant output from all feature maps without further input. This is followed by a series of 1x1 convolutions that reduce the feature dimensions of each pixel to a raw RGB value and then output a rectified color per pixel via a sigmoid activation layer.
III-E Training Details
Due to the use of an image decoder, rays must be processed in patches rather than sampled individually throughout the image. This leads to a potential complication from edge effects within the image decoder because of the downsampling nature of convolutional layers and the padding that may compensate for it. Because we are rendering an image based on patches and not in its entirety due to memory constraints, a simple padding scheme should not be used when training or during inference. If using padding in the convolutional layers, discontinuities will appear during training and inference that affect the quality of the final images. In order to adjust for this, we over-sample our patches and do not utilize padding within the convolutional layers of the image decoder. This over-sampling effectively serves as padding for the desired final shape after decoding except that actual grid values are used rather than approximations such as zero-padding or reflectance. In this case the over-sampling must be tailored to the architecture of the image decoder itself such that the final dimensions of the output image are appropriate. Given our architecture, we over-sample by eight ”pixels” in each spatial dimension for each feature map.
We use a NVIDIA RTX 3090 GPU for training with a learning rate of $51e-$ 5 for both the feature grids as well as the image decoder. The training is run for 500,000 iterations with a patch from one image per training iteration. Patches are sized so as to be an integer divisible by the width and height of the actual image while also being evenly divisible by two for each scale. In our case we chose a patch size of 360x640 for every experiment except Okutama video 1.1.1, in which case we chose a smaller patch size of 64x72. In most cases the larger patch size showed considerable performance improvements over using smaller patch sizes III.
@@ IV-A1 Okutama-Action Dataset
TABLE I: PSNR Comparison on Okutama-Action and UG2 Validation Subsets
Okutama K-Planes Extended K-Planes TK-Planes
1.1.1 18.23 32.61 31.17
1.2.2 14.45 32.63 32.93
1.2.6 14.37 27.89 28.97
UG2 K-Planes Extended K-Planes TK-Planes
UAV-133 17.11 27.56 27.88
Okutama K-Planes Extended K-Planes TK-Planes
1.1.1 20.55 24.07 27.36
1.2.2 15.63 24.52 26.47
1.2.6 16.31 20.81 21.76
UG2 K-Planes Extended K-Planes TK-Planes
UAV-133 15.54 17.51 16.91
TABLE III: Dynamic PSNR for Small and Large Patch-Size on Okutama-Action and UG2 Validation Subsets for TK-Planes
Okutama 64x72 360x640
1.1.1 27.36 25.20
1.2.2 22.59 26.47
1.2.6 18.40 21.76
UG2 64x72 360x640
UAV-133 15.38 16.91
(a) Okutama
(b) K-Planes
(c) Extended
(d) TK-Planes
Figure 3: A comparison of dynamic regions from a subsection of video 1.1.1 in Okutama-Action. a.) Okutama ground truth b.) stock K-Planes c.) Extended K-Planes d.) TK-Planes. We highlight the improved rendering quality generated by TK-Planes over the other methods on these challenging dynamic scenes.
(a) Okutama
(b) K-Planes
(c) Extended
(d) TK-Planes
Figure 4: A comparison of dynamic regions from a subsection of video 1.2.2 in Okutama-Action. a.) Okutama ground truth b.) stock K-Planes c.) Extended K-Planes d.) TK-Planes. We highlight the improved rendering quality generated by TK-Planes over the other methods on these challenging dynamic scenes.
(a) Okutama
(b) K-Planes
(c) Extended
(d) TK-Planes
Figure 5: A comparison of dynamic regions from a subsection of video 1.2.6 in Okutama-Action. a.) Okutama ground truth b.) stock K-Planes c.) Extended K-Planes d.) TK-Planes. We highlight the improved rendering quality generated by TK-Planes over the other methods on these challenging dynamic scenes.
(a) UG2
(b) K-Planes
(c) Extended
(d) TK-Planes
Figure 6: A comparison of dynamic regions from a subsection of video UAV-133 in UG2. a.) UG2 ground truth b.) stock K-Planes c.) Extended K-Planes d.) TK-Planes. The UG2 dataset is particularly challenging, extended K-Planes performs better in some cases while TK-Planes in others.
The Okutama-Action dataset [3] is collected using unmanned aerial vehicles (UAVs). It comprises scenes captured at various altitudes, angles, and time periods, containing different humans in 12 different action classes. This dataset presents challenges due to dynamic transitions between actions, humans of multiple concurrent actions, and actors labeled with multiple actions. We train on the three different scenes: 1.1.1, 1.2.2, 1.2.6. Each of these scenes offers distinct views of a baseball diamond at different times of day with varying numbers of people present, see Figures 7(a), 7(b), 7(c).
@@ IV-A2 UG2 Dataset
We also sample a video from the UG2 dataset [39] which consists of a large variety of videos from varying styles of aerial vehicles, including piloted aircraft and UAVs such as fixed wing gliders and quadrocopters. Although many of the videos within UG2 are not applicable for this given application, video 133 within the test set for UAV videos offers a challenging view of a cricket match in progress. This particular video contrasts well with the scenes from Okutama-Action, offering smaller and more dynamic targets moving with less precise and scripted actions 7(d).
IV-B Analysis
Because we are focused on dynamic portions of a scene, we calculate not only overall PSNR values but also what we call Dynamic-PSNR values which represent the average PSNR between each bounding box of people within each frame. With respect to tasks such as object detection, pose recognition and action recognition, the fidelity of the target object is key and thus the most pertinent for validation. As can be seen from Tables I and II, the overall performance of TK-Planes beats stock and extended K-Planes in most experiments. Furthermore, Table III shows how important hyperparamter tuning is with respect to TK-Planes. In all but the first experiment, larger patch-sizes provides upwards of 3.88 points of improvement for Dynamic-PSNR. Given that TK-Planes has more hyperparameters to adjust than stock or extended K-Planes, there is a larger space for exploration in achieving improved results which indicates that TK-Planes may further improve PSNR values compared to stock and extended K-Planes.
The ability to utilize an image decoder and to store more abstract feature vectors within the grids allows for the network to learn how to render a person versus rendering individual pixel values. In the latter case such as seen in Extended K-Planes, a higher level of noise is captured in the renderings while also losing some of the detail and sharpness of smaller portions of the people. Figures 3(d), 4(d), and 5(d) in particular shows enhanced detail from TK-Planes with respect to limbs and clothing detail. One of the main benefits from this is in rendering synthetic data for the previously mentioned tasks that require specific limb placement and outlines, such as action recognition.
Samples from the Okutama experiments and Tables I and II shows that TK-Planes can ultimately render higher fidelity imagery when compared to Extended K-Planes, and especially when compared to standard K-Planes which fails to render details in many cases. However, the final experiment from UG2 of the UAV cricket scene shows more mixed results. The overal PSNR value for TK-Planes is higher when compared to stock and extended K-Planes, but the Dynamic-PSNR value lower when compared to extended K-Planes. This scene is particularly difficult with small dynamic portions and large unbounded backgrounds. Figure 6 compares renderings of people from all three methods with that of the ground truth image from UG2. Although the background is largely captured, the dynamic portions of the scene are very small resulting in a lack of detail with respect to the people in the scene. The top two rows show extended K-Planes rendering more accurate people while the bottom two rows show TK-Planes renderings to be more accurate. Due to the difficulty of this scene, none of the methods perform particularly well, but TK-Planes has shown that hyperparameter tuning can play a large role in its overall performance which leads to the intuition that TK-Planes is not optimized for this particular scene and may yet achieve better performance.
IV-C Failure Modes and Limitations
In order to explain the failure cases such as seen in Figure 6(d), there are a variety of factors that contribute to this dichotomy between Extended and TK-Planes. First, the training method for TK-Planes has a larger number of hyperparameters, the most likely of which to affect specific frames having portions of lower fidelity being the sampling rate of dynamic portions of the scene and the overall patch size per image during training. Second, each scene for Okutama and UG2 does not provide ground truth camera poses which are a necessity for many NeRF algorithms. To compensate, we run advanced structure from motion algorithms [19] to estimate the camera poses for each image within a scene. It is possible that these estimated poses are large enough confounding factors between scenes with respect to the overall PSNR of dynamic portions of a scene.
A further limitation that is also seen in Extended K-Planes is that of limited novelty beyond training poses and time-slices. In Extended K-Planes, the further from the training poses and time-slices a rendering pose and time departed, the more distorted the rendered result would be. A similar limitation is also present within TK-Planes. This is ultimately due to the difficult restrictions of the UAV datasets that have tight correlations between poses and time-slices, as well as limited diversity in overall poses at varying time-slices.

### v2

@@
Correspondence to cmaxey@umd.edu
In this paper, we present a new approach to bridge the domain gap between synthetic and real-world data for unmanned aerial vehicle (UAV)-based perception. Our formulation is designed for dynamic scenes, consisting of small moving objects or human actions. We propose an extension of K-Planes Neural Radiance Field (NeRF), wherein our algorithm stores a set of tiered feature vectors. The tiered feature vectors are generated to effectively model conceptual information about a scene as well as an image decoder that transforms output feature maps into RGB images. Our technique leverages the information amongst both static and dynamic objects within a scene and is able to capture salient scene attributes of high altitude videos. We evaluate its performance on challenging datasets, including Okutama Action and UG2, and observe considerable improvement in accuracy over state of the art neural rendering methods.
@@ I Introduction
Figure 1: Example neural rendering results. Our target dynamic scene contains small, fast-moving people (in red box) captured using only a monocular camera. Despite this challenging scenario, our proposed method produces high-quality renderings for dynamic scene.
@@
We have evaluated the performance on some challenging UAV datasets, including Okutama Action [3] and UG2 [40], which consist of high altitude videos. We show that TK-planes can generate high fidelity imagery of UAV scenes, capturing small dynamic details well and outperforming previous state of the art neural rendering methods.
  • Development of a novel grid-based NeRF architecture that utilizes multiple scales of feature space in order to render frames from a dynamic UAV scene.
  • Extension of previous pipelines for training grid-based NeRF architectures to include an image decoder.
  • Showcasing high fidelity renderings with PSNR and dynamic-PSNR (DPSNR) values exceeding previous methods such as K-Planes [9], Extended K-Planes [21] and 4D Gaussian Splatting [43].
@@ II Related Works
Neural Rendering for Dynamic Scenes. To represent 3D dynamic scenes, various NeRF research explores 4D neural radiance fields representation with an additional temporal dimension [6, 45, 10, 41]. Instead of embedding continuous time dimension into the radiance fields, some of variants uses time-varying latent code [18, 27]. While these 4D NeRF show capability to represent complicated dynamic scenes, they generally suffer from inherent representational redundancy when they learn 4D contents from 2D imagery only. The other major variants of NeRF incorporates canonical scene and a deformation field [26, 30]. However, these methods demonstrate limited capability to learn topological changes or large motion in the scene due to their heavy dependence on the canonical scene. Recently, Gaussian Splatting [16] has become a state of the art research topic for neural rendering, producing accurate results with fast rendering speeds. [43] applies Gaussian Splatting to temporal scenes, outperforming previous 4D neural rendering methods in both speed and accuracy.
@@ III-A Novel View Synthesis for Dynamic Scenes
Given a difficult domain such as UAV video with limited available training data and a small dynamic to static pixel ratio, our approach utilizes NeRF in feature space in order to better capture and render objects of interest within a scene, e.g. dynamic objects. We employ a grid-based Neural Radiance Field, the grids of which are similar to those from K-Planes [9]. A key difference being that the grids in our algorithm store feature vectors that do not directly encode for RGB values but rather for more abstract high level concepts within the scene. As such, our NeRF algorithm outputs a series of feature vectors and density values for a given ray which are then combined via typical volumetric rendering. These combined feature vectors form a feature map as opposed to an RGB image. The output feature maps are decoded by a convolutional neural network in order to generate a final rendered RGB image.
III-B Neural Radiance Field Overview
To understand our algorithm in more detail, a brief overview of Neural Radiance Fields follows. Neural radiance fields are a type of volumetric rendering that utilize neural networks to output a color value and density given an input camera pose. From a designated camera pose, rays are cast into a scene and incremental points along each ray are sampled. Typically, two Multi-layer Perceptrons (MLPs) are employed, one for spatial density and another for a directionally dependent RGB value. The spatial density MLP takes as input a position within 3D space and outputs a density value as well as a feature vector. This feature vector and the given ray direction are fed into the second MLP which outputs a RGB color value. The density and color value at each sampled point along the ray (Eq. 3-(1)) are then accumulated via volumetric rendering. Eq. 3-(2) depicts the overall color value given a ray wherein the point specific color value, density, and transmittance (Eq. 3)-(3) up to the given point are multiplied and integrated. The transmittance is a representation of how much light will pass through all previous points and thus how much the current point contributes to the overall color value. This version of NeRF relies on the MLPs learning an implicit 3D structure of a scene.
$r(t)$ $=o+td$ (1)
[…]
Alternative to an implicit method, some NeRF algorithms implement an explicit representation of a scene stored in a data structure such as a matrix [15, 39] or a series of grids [9, 4]. These data structures contain feature vectors that learn details of a scene and when passed through a density and color MLP, output density and color values respectively. The explicit spatial, and temporal if applicable, layout of the data structures offloads the necessity for implicitly learning the spatial and temporal structure of the scene within the MLPs. The benefit of this is that smaller MLPs can be used resulting in quicker training and inference times. The downside is increased memory usage, in particular for matrix representations, though the use of factored grid representations [9, 4] helps to alleviate this requirement.
@@ III-C Tiered Planes in Feature Space
Our main approach operates in a conceptual feature space rather than an RGB space with respect to the output of the NeRF algorithm. The main purpose of focusing on feature space rather than RGB space is to learn feature vectors that represent cohesive objects of a scene, such as people, as opposed to learning solely a color and density value. This helps to alleviate the issue of poor fidelity of high interest objects across different poses and time-slices. We use the same grid representation as seen by K-Planes [9] with changes influenced from the extended version as seen in [21]. Figure 2 shows our system in full which includes abstract feature grids (up to $n$) as well as an image decoder. Our NeRF model outputs a feature map rather than a rgb image and the image decoder processes the feature map into a final image.
An ancillary benefit of TK-Planes is a vast decrease in training and inference times when compared to previous state of the art NeRF methods. In general, explicit factored representations of a scene improve the speed of training and inference; however, a bottleneck still exists due to the ray casting and ray sampling required for NeRF techniques, wherein a large number of poses (on the order of 10’s of thousands) within a scene are processed. A reduction in the number of points required during sampling has a direct impact on the speed of training and inference. Our method allows for a considerable reduction in the number of rays needed to render an image as well as the number of points along each ray that need to be sampled. This is due to the fact that the output of our NeRF model is a feature map rather than a RGB image and said feature maps can have height and width dimensions much smaller than that of the final image. For instance, the final image dimensions in the Okutama-Action dataset are 720x1280 yet the feature map dimensions used in our experiments for this dataset start at 45x80, an overall reduction factor of 256 with respect to the number of rays needed to render a single image.
III-D TK-Planes Architecture Details
As seen in [21], we add the additional planes for spatial features that are dynamic. The resulting static and dynamic features are described as follows:
@@ III-D TK-Planes Architecture Details
Figure 2: Overview of our tiered planes algorithm: An arbitrary $n$ number of grids are shown, representing different scales of feature space, with larger scales (earlier feature maps) capturing more abstract scene details. Note that the algorithm uses nine grids in total per scale, three static spatial, three dynamic spatial and three dynamic spatio-temporal. Feature vectors are processed via concurrent MLPs to output a density value and final feature vector. These are accumulated with volumetric rendering into a feature map. Feature maps from each set of grids are input into corresponding stages of the image decoder.
wherein $f_{s}$ is a function of the static spatial planes and $f_{d}$ is a function of the dynamic spatial and temporal planes and $\prod$ represents the Hadamard product. In our version, each feature vector $f_{*}(\mathbf{q})_{c}$ is a $D$ dimensional vector wherein $D$ is a hyperparameter. From 5, a grouping of feature vectors $f(\mathbf{q})$ is calculated by concatenating $f_{s}$ and $f_{d}$ in the batch dimension such that each spatial and dynamic feature vector can be processed independently. $f(\mathbf{q})$ is then input into the MLP networks to output a density value and another feature vector for each static and dynamic feature vector. Volumetric rendering along the ray returns a final spatial and dynamic feature vector value per pixel of an output feature map at a given feature space scale, $d_{s}$. To maintain a balance between speed, accuracy and memory usage, our experiments include using up to two feature space tiers which allows for training using an entire image rather than an image patch.
K-Planes [9] and its extended version [21] use multi resolution grids offering different levels of scene details. [43] also uses a k-planes style multi-resolution grid system to sample small feature vectors for processing gaussian parameters. The resulting feature vector from each level can either be summed or concatenated before being processed by the MLPs and in the case of NeRF models, then accumulated via volumetric rendering. In these cases, the volumetric rendering is the final step. However, in our algorithm it is more intuitive to separate the different levels of detail as input into appropriate stages of the image decoder rather than as input entirely into the front of the decoder. Convolutional image decoders operate on high-level abstract features to low-level detail features as their stages progress and typically benefit from feature map input at varying stages [33]. This allows the grids at varying scales of feature space to learn appropriate details with respect to rendering an output image.
@@
[…]
III-E Image Decoder
[…]
After all upsampling stages, a final stage combines the static and dynamic feature maps into an output image. At this stage, the static and feature maps that were processed independently in previous stages are concatenated along their channel dimensions.
III-F Training Details
Due to the use of an image decoder, rays must be processed in patches rather than sampled randomly throughout the image. In the case of using only one or two tier scales, the entire image can be used in one training pass as compared to a subset of image pixels in previous NeRF methods. This benefits the quality of the final rendering and contributes to the large decrease in rendering times.
In order to get estimated camera poses, we use pix-sfm [19] with feature type superpoint, matcher type superglue and sfm tool hloc. This is applied to all images from each data subset before splitting into training and validation. For calculating dynamic PSNR, bounding boxes are needed for the dynamic regions of a scene, e.g. the people. Okutama-Action has ground truth bounding boxes, however UG2 does not. In order to rectify this, we used YOLOv8 to extract bounding boxes for each person in the UG2 video.
We use a NVIDIA RTX 3090 GPU for training with a learning rate of $31e-$ 3 for both the feature grids as well as the image decoder. The training is run for 100,000 iterations using an entire image per training iteration. Hyperparameter details between each method are kept as close as possible so as to offer a fair comparison between each method. Although 4D Gaussian Splatting uses a different mechanic than NeRF, it does use a similar grid system with multiple scales as K-Planes. The feature vector sizes within these grids are set so as to equal that of two tiers of the feature vectors in TK-Planes.
@@ IV-A1 Okutama-Action Dataset
The Okutama-Action dataset [3] is collected using unmanned aerial vehicles (UAVs). It comprises scenes captured at various altitudes, angles, and time periods, containing different humans in 12 different action classes. This dataset presents challenges due to dynamic transitions between actions, humans of multiple concurrent actions, and actors labeled with multiple actions. We train on the three different scenes: 1.1.1, 1.2.2, 1.2.6. Each of these scenes offers distinct views of a baseball diamond at different times of day with varying numbers of people present, see Figures 4(a), 4(b), 4(c).
@@ IV-A2 UG2 Dataset
We also sample a video from the UG2 dataset [40] which consists of a large variety of videos from varying styles of aerial vehicles, including piloted aircraft and UAVs such as fixed wing gliders and quadrocopters. Although many of the videos within UG2 are not applicable for this given application, video 133 within the test set for UAV videos offers a challenging view of a cricket match in progress. This particular video contrasts well with the scenes from Okutama-Action, offering smaller and more dynamic targets moving with less precise and scripted actions 4(d).
@@
TABLE I: PSNR Comparison on Okutama and UG2 Validation Subsets
Okutama K-Planes [9] Ext. K-Planes [21] 4D Gaussian [43] TK-Planes (Ours)
1.1.1 18.23 32.61 33.95 34.41
1.2.2 14.45 32.63 34.45 34.40
1.2.6 14.37 27.89 29.35 30.28
UG2 K-Planes [9] Ext. K-Planes [21] 4D Gaussian [43] TK-Planes (Ours)
UAV-133 17.11 27.56 30.47 31.52
Figure 5: A comparison of dynamic regions from a subsection of video 1.1.1 in Okutama-Action. a.) Okutama ground truth, b.) stock K-Planes, c.) Extended K-Planes, d.) 4D-Gaussian, e.) TK-Planes. We highlight the improved rendering quality generated by TK-Planes over the other methods on these challenging dynamic scenes.
Okutama K-Planes [9] Ext. K-Planes [21] 4D Gaussian [43] TK-Planes (Ours)
1.1.1 20.55 24.07 24.56 28.17
1.2.2 15.63 24.52 24.69 27.97
1.2.6 16.31 20.81 20.66 23.92
Okutama K-Planes [9] Ext. K-Planes [21] 4D Gaussian [43] TK-Planes (Ours)
UAV-133 15.54 17.51 16.44 21.02
Figure 6: A comparison of dynamic regions from a subsection of video 1.2.2 in Okutama-Action. a.) Okutama ground truth, b.) stock K-Planes, c.) Extended K-Planes, d.) 4D-Gaussian, e.) TK-Planes. We highlight the improved rendering quality generated by TK-Planes over the other methods on these challenging dynamic scenes.
Figure 7: A comparison of dynamic regions from a subsection of video 1.2.6 in Okutama-Action. a.) Okutama ground truth, b.) stock K-Planes, c.) Extended K-Planes, d.) 4D-Gaussian, e.) TK-Planes. We highlight the improved rendering quality generated by TK-Planes over the other methods on these challenging dynamic scenes.
Figure 8: A comparison of dynamic regions from a subsection of video UAV-133 in UG2. a.) UG2 ground truth, b.) stock K-Planes, c.) Extended K-Planes, d.) 4D-Gaussian, e.) TK-Planes. The UG2 dataset is particularly challenging; however, it can be seen that TK-Planes performs admirably compared to the other methods.
IV-B Analysis
Because we are focused on dynamic portions of a scene, we calculate not only overall PSNR values but also what we call Dynamic-PSNR (DPSNR) values which represent the average PSNR of each bounding box for people within each frame. With respect to tasks such as object detection, pose recognition and action recognition, the fidelity of the target object is key and thus the most pertinent for validation. As can be seen from Tables I and II, the overall performance of TK-Planes outperforms stock and extended K-Planes in all experiments, up to 1.8 PSNR in Okutama tests and 3.96 PSNR in the UG2 test. With respect to DPSNR, TK-Planes takes a wider lead with up to 4.1 DPSNR improvement in Okutama tests and 3.51 DPSNR in the very difficult UG2 test.
Furthermore, TK-Planes with two tiers of feature vector grids runs 10x faster when rendering than previous versions of K-Planes. A one tier version of TK-Planes still outperforms previous K-Planes methods while running 25x faster when rendering. This is due to the lower number of rays and sampling points along each ray required by TK-Planes versus K-Planes or other previous state of the art NeRF methods. It is noted that the version of TK-Planes in this paper uses the simple ray sampling method of uniform sampling, compared to the more advanced proposal sampling method of K-Planes and extended K-Planes. See [22] for more comparisons between one and two tier TK-Planes.
4D Gaussian Splatting [43] is a state of the art neural rendering method that outperforms many previous state of the art NeRF methods, including K-Planes. With respect to UAV videos with estimated camera poses, TK-Planes outperforms 4D Gaussians in the Okutama tests by up to 0.93 PSNR, and 1.05 PSNR on the difficult UG2 test. The performance gap is widened with respect to DPSNR, with up to 3.61 DPSNR improvement in the Okutama tests and 4.58 DPSNR in the UG2 test. X5, 6, 7 and 8 showcase some examples of DSPNR regions, e.g. parts of the frame with people. The performance improvement for dynamic people can be qualitatively seen in these images, in particular for the UG2 test, in Figure 8. See [22] for more comparisons between K-Planes, extended K-Planes, TK-Planes and 4D Gaussians.